OpenStudy (anonymous):

Calculate the discriminant. x2 – 2x + 4 = 0 A. –12 B. 12 C. 20 D. 0

hero (hero):

\(D = b^2 - 4ac\)

hero (hero):

If given a quadratic equation of the form \(ax^2 + bx + c = 0\), then the value of the discriminant can be found using \(D = b^2 - 4ac\)

hero (hero):

In this case, a = 1 b = -2 c = 4

hero (hero):

Compute this: \((-2)^2 - 4(1)(4)\)